From narrow range pH values. Use [H^+] = 10^-pH;
Answer #1

For H2O (unboiled):

Narrow pH range = 5.5

[H+] = 10-5.5 = 3.16 x 10-6 M

[OH-] = 10-14 / [H+] = 10-14 / 3.16 x 10-6 = 3.16 x 10-9 M

The complete table is shown below for all solutions following the same procedure:

pH WideNarrow range| range Acidic/Neutral Solution [H, M[OH], M /Basic H20 (unboiled4 H20 (boiled) Nacl NaAc NH4CI ZnC12 KAl(